OpenStudy (buggiethebug):

Determine if each of the following statements is always true, sometimes true, or never true. Explain a) If Lim(x-->0)f(x)=0, then f(x) is continuous at x=0

OpenStudy (anonymous):

Sometimes true. Consider \[f(x)=\frac{x^2}{x}\] There's a hole at x = 0, so f(x) is not continuous at x = 0, despite the fact that \[\lim_{x\to0}f(x)=0.\]

OpenStudy (buggiethebug):

and what about this "If g(1)=3, then lim(x-->1)g(x)=3

OpenStudy (anonymous):

Also sometimes true. Take the following function: \[g(x)=\begin{cases}x&\text{if $x\not=1$}\\ 3&\text{if $x=1$}\end{cases}\] By this definition of g(x), you have g(1) = 3. However, \[\lim_{x\to1}g(x)=\lim_{x\to1}x=1.\]